# Some generalizations of Camina Pairs and orders of elements in cosets

### v1

@@ 1. Introduction
Our first theorem establishes the equivalences of the $\mathrm{(CI)}$ and $\mathrm{(F)}$ conditions for any finite group $G$ and nontrivial proper subgroup $H$.
@@
Let $G$ be a finite group and let $H$ be a nontrivial proper subgroup of $G$. Let $N$ be the normal closure of $H$ in $G$. If the pair $(G,H)$ satisfies condition $\mathrm{(F^{\prime})}$, then $(G,N)$ satisfies $\mathrm{(F^{\prime})}$ and $N$ is nilpotent.
@@
Let $G$ be a finite group and let $H$ be a nontrivial proper subgroup of $G$. Study the structure of $G$ if the pair $(G,H)$ satisfies the condition: for every $x\in G\setminus H$, the elements in the coset $xH$ all have the same order as $x$.
The main inspiration for our definition of condition $\mathrm{(F)}$ follows from work of Navarro and Guralnick in [GN]. In that paper, the authors investigate the structure of finite group $G$ with a conjugacy class $K$ such that $K^{2}$ is a conjugacy class of $G$. This is a special case of a conjecture proposed by Arad, Stavi and Herzog in [ASH] stating that in a finite nonabelian simple group, the product of two nontrivial conjugacy classes cannot be a single class. In [GN], they proved that if $K^{2}$ is a conjugacy class of $G$, then $N:=[x,G]$ is solvable, where $x\in K$. Moreover, all elements of the coset $xN$ are conjugate in $G$. Inspired by this, they also show that if $N$ is a normal subgroup of $G$ and $x\in G$ and that either all elements of $xN$ are $G$-conjugate or all elements of $xN$ have odd order, then $N$ is solvable. As before, dropping the normality assumption, we can prove the following result.
@@ 3. $(c)$
This result may be known to experts. In fact, it has been shown that if $x\in G$ and $xy$ is $p$-regular for every $p$-regular element $y\in G$, then $x\in O_{p^{\prime}}(G)$. This was proved independently by Robinson [R] and Flavell [F]. Robinson’s proof is character-theoretic, while Flavell’s argument is group-theoretic. Both proofs do not use the classification of finite simple groups. Our proof of Corollary 5 is also classification-free. Corollary 5 can be considered a dual of Theorem D in [BLM+], where it is shown that if $x\in G$ is $p$-regular such that the order of $xy$ is divisible by $p$ for all nontrivial $p$-elements $y\in G$, then $x\in O_{p^{\prime}}(G)$. This is proved modulo some conjecture on finite almost simple groups.
@@ 2. Preliminaries
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(F)}$ and $G$ is a $p$-group for some prime $p$. Then $H$ is a normal subgroup of $G$.
@@
Since $G$ is a finite $p$-group, it is nilpotent and possesses an upper central series
@@
Because $Z_{1}=\mathbf{Z}(G)$ is a nontrivial normal $p$-subgroup of $G$, by Lemma 2.6, $Z_{1}\leqslant H$. If $Z_{1}=H$, then the claim follows. Suppose instead that $Z_{1}<H$. Then, by Lemma 2.5, the pair $\left(G/Z_{1},H/Z_{1}\right)$ satisfies condition $\mathrm{(F)}$. Applying Lemma 2.6 to the pair $(G/Z_{1},H/Z_{1})$ yields $Z_{2}/Z_{1}=\mathbf{Z}(G/Z_{1})\leqslant H/Z_{1}$ and hence $Z_{2}\leqslant H$. If $Z_{2}=H$, then we are done. Otherwise, we continue the process: the pair $(G/Z_{2},H/Z_{2})$ satisfies condition $\mathrm{(F)}$ so $Z_{3}/Z_{2}=\mathbf{Z}(G/Z_{2})\leqslant H/Z_{2}$, and thus $Z_{3}\leqslant H$. By repeating this argument, we either eventually reach $H=Z_{i}$ for some $i$, in which case $H\unlhd G$ or else, we find that $Z_{k-1}\leqslant H$. Since $G^{\prime}\leqslant Z_{k-1}\leqslant H$, $H$ is normal in $G$ as well. ∎
@@
Thus, $Kh^{g}=K$ for all $g\in G,h\in H$. Since every $n\in N$ is the product of finitely many such conjugates $h^{g}$, it follows that $Kn=K$ for all $n\in N$, so $KN=K$. Therefore, $xN\subseteq K=x^{G}$ for every $x\in G\setminus N$, and hence $(G,N)$ is a Camina pair. ∎
@@
Since $N$ is nilpotent by Lemma 2.10, its derived subgroup $N^{\prime}$ is a proper characteristic subgroup of $N$. Since $N$ is normal in $G$, $N^{\prime}\unlhd G$. Moreover, since $N$ is the smallest normal subgroup of $G$ containing $H$ and $H$ is not normal in $G$, we must have $N^{\prime}<H$. Therefore, $H\unlhd N$ and $N/H$ is abelian. ∎
@@ 3. Proof of Theorem 1
where $\theta_{1},\dots,\theta_{t}$ are the distinct irreducible characters of $N$ that are conjugate under the action of $G$. Since $\chi_{H},1_{H}]\neq 0$, we may assume without loss of generality that the trivial character $1_{H}$ lies under $\theta_{1},\ldots,\theta_{s}$ for some integer $s$ with $1\leqslant s\leqslant t$. By Lemma 2.4, we know that $H\unlhd N$. Applying Clifford’s Theorem again to each $\theta_{i}$ with $1\leqslant i\leqslant s$, the restrictions of all these characters $\theta_{1},\ldots,\theta_{s}$ to $H$ are multiple of $1_{H}$. Consequently, for $1\leqslant i\leqslant s$, $\theta_{i}$ is an irreducible character of the quotient group $N/H$, which is abelian by Lemma 2.4. Hence these $\theta_{i}$ are all linear characters of $N$. Finally, since $\theta_{1},\dots,\theta_{t}$ are conjugate in $G$, they all have the same degree. Therefore, all $\theta_{i}$, for $1\leqslant i\leqslant t$, are linear characters, as required.
@@
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(CI)}$ and that $H$ is not normal in $G$. Let $N$ be the normal closure of $H$ in $G$. Assume $\chi\in\mathrm{Irr}(G\mid N)$ such that $\chi_{H},1_{H}]=0$. Then $\chi(xh)=\chi(x)$ for all $x\in N\setminus H$ and $h\in H$.
@@
where the $\theta_{i}$ for $i=1,2,\ldots,t$ are all the distinct irreducible characters of $N$, which are conjugate in $G$. Since $\chi_{H},1_{H}]=0$, we can find a nontrivial irreducible constituent $\mu_{1}$ of $\chi_{H}$. We may assume that $\theta_{1},\ldots,\theta_{l_{1}}$, for some $l_{1}\leqslant t$, are precisely the irreducible constituents of $\chi_{N}$ lying over $\mu_{1}$. Analogous to the argument in Proposition 3.1, we claim
@@
Let $N$ denote the normal closure of $H$ in $G$. By Lemma 2.2, $(G,N)$ is a Camina pair. That is, for every $\in G\setminus N, xx$ is conjugate to $xn$ for all $n\in N$. In particular, for all $x\in G\setminus N$ and $h\in H\subseteq N$, $x$ is conjugate to $xh$, as required for condition $\mathrm{(F)}$. It remains to consider the case when $x\in N\setminus H$. We claim that for all $\in N\setminus H, xx$ is conjugate to $xh$ for all $h\in H$. This is equivalent to for all $x\in N\setminus H$, $\chi(x)=\chi(xh)$ for all $h\in H$ and $\chi\in\mathrm{Irr}(G)$. Indeed, for any $\chi\in\mathrm{Irr}(G/N)$, we have $\chi(x)=\chi(xh)=\chi(1)$ for all $x\in N,h\in H$ as $N$ is contained in the kernel of $\chi$. For $\chi\in\mathrm{Irr}(G\mid N)$, the equality $\chi(x)=\chi(xh)$ for all $x\in N\setminus H$, $h\in H$ followed from Propositions 3.1 and 3.2. Thus, all irreducible characters of $G$ are constant on the cosets $xH$ for $x\in N\setminus H$, so $xH\subseteq x^{G}$ for all $x\in N\setminus H.$ Hence $(G,H)$ satisfies condition $\mathrm{(F)}$, completing the proof. ∎
@@
This follows from Propositions 3.3 and 3.4. ∎
@@ 4. Proof of Theorem 2
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(F^{\prime})}$ and that $H$ is not normal in $G$. Let $N$ be the normal closure of $H$ in $G$. Then $N$ is solvable and has a normal $p$-complement for some prime $p$.
@@
Since $H$ is a proper subgroup of $G$, $\Delta_{H}(G)=G\setminus\cup_{g\in G}H^{g}$ is non-empty. By Corollary 4.4, we have $\Delta_{H}(G)=G\setminus N$ and so $N<G$. Let $x\in\Delta_{H}(G)$ and let $K=x^{G}$. By Lemma 4.3, $KN\subseteq K\cup K^{-1}$, then $K\cup K^{-1}$ is a union of cosets of $N$. By [B, Theorem 3.3(a)], $N$ is solvable. Since $G/N$ is nontrivial, we can find $x\in G\setminus N$ such that $x$ is a $p$-element for some prime $p$. Then, by [B, Theorem 3.3(c)], $N$ has a normal $p$-complement. ∎
@@
We now prove that $G$ is a Frobenius group with Frobenius kernel $N$, so that $N$ is nilpotent by Thompson’s Theorem, yielding a contradiction. Recall that $N$ is a non-nilpotent normal $p$-complement of $G$. Since $G=PN$, with $N$ a normal $p$-complement and $P\cap N=1$, it suffices (see [I, Problem 7.1]) to show that $\mathbf{C}_{G}(n)\subseteq N$ for every $1\neq n\in N.$ Suppose, for contradiction, that there exists $x\in\mathbf{C}_{G}(n)\setminus N$. Then $x$ is a $p$-element and $xn=nx$. Since $n$ is a $p^{\prime}$-element, the order of $xn$ is $o(x)o(n)$, so $xn$ is not a $p$-element. But since $x\in G\setminus N$, condition $\mathrm{(F^{\prime})}$ implies that $xn\in xN\subseteq x^{G}\cup(x^{-1})^{G}$, a set consisting of only $p$-elements. This contradiction shows that $\mathbf{C}_{G}(n)\subseteq N$, and thus $G$ is a Frobenius group with Frobenius kernel $N$. By Thompson’s theorem [Th, Theorem 1], $N$ is nilpotent, contradicting our assumption. Therefore, $N$ must be nilpotent, completing the proof. ∎
@@ 2. $(2)$
Let $K_{0}$ be a subgroup of $G$ generated by all $p$-regular elements of $G$. Then $K_{0}\unlhd G$ because the set of $p$-regular elements is closed under conjugation. We will show that $K_{0}=K$, and so $K\unlhd G$. From the previous paragraph, $K_{0}\subseteq K$. Now consider an arbitrary element $hK_{0}\in H/K_{0}$, and write $h=h_{p}h_{p^{\prime}}$ where $h_{p},h_{p^{\prime}}$ are the $p$-part and $p^{\prime}$-part of $h$, respectively. Then $h_{p^{\prime}}\in K_{0}$, so $hK_{0}=h_{p}K_{0}$ has $p$-power order in $H/K_{0}$. Thus all elements of $H/K_{0}$ are $p$-elements, and so $H/K_{0}$ is a $p$-group. Hence $K\subseteq K_{0}$, and we conclude that $K=K_{0}\unlhd G$. Moreover, $|G:H|$ is a $p$-power and $H/K$ is a $p$-group, which implies that $G/K$ is a $p$-group.
@@
The following lemma is obvious.
Lemma 5.3.
@@
Proof.
Since $K\lhd N$, $K\cap S_{i}\lhd S_{i}$. Moreover, since $S_{i}\cong S$ is nonabelian simple for all $1\leqslant i\leqslant k$, either $K\cap S_{i}=1$ or $K\cap S_{i}=S_{i}$. Without loss of generality, we can assume that $K\cap S_{i}=S_{i}$ for all $1\leqslant i\leqslant s$ and $K\cap S_{i}=1$ for all $s+1\leqslant i\leqslant k$ with $s<k$. It follows that $S_{i}\leqslant K$ for all $1\leqslant i\leqslant s$ and so $S_{1}\times\ldots\times S_{s}\leqslant K$. We will show that $S_{1}\times\ldots\times S_{s}=K$. If $x\in K\setminus S_{1}\times\ldots\times S_{s}$, then there must exist $m\in\left\{s+1,\ldots,k\right\}$ such that the projection $p_{m}:N\rightarrow S_{m}$ maps $x$ to a nontrivial element of $S_{m}$. Hence, $K$ intersects $S_{m}$ nontrivially for some $s+1\leqslant m\leqslant k$, a contradiction. Therefore, $K=S_{1}\times\ldots\times S_{s}$ for some $s<k$. Consequently, $N/K\cong S_{s+1}\times\ldots\times S_{k}$ is not solvable. ∎
@@
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(O)}$. Then either $O^{2}(H)\unlhd G$ or $H$ is solvable.
@@
Let $g\in G\setminus H$ be an element of even order. Suppose, for contradiction, that the coset $gH$ contains an element of odd order. Then there exists $h\in H$ such that $x=gh\in gH$ has odd order. Since $x\in G\setminus H$ and has odd order, the pair $(G,H)$ satisfying condition $\mathrm{(O)}$ implies that every element in the coset $xH$ also has odd order.
Now, observe that $h^{-1}\in H$, so $xh^{-1}=(gh)h^{-1}=g\in xH$. Thus, $g$ must also have odd order, contradicting our assumption that $g$ has even order. Therefore, all elements in $gH$ must have even order.
@@
Suppose, for contradiction, that $N=H_{G}>1$. Then $N\trianglelefteq G$, $N\leqslant H$, and $1<N<H$. By Claim 3, the pair $(G/N,H/N)$ satisfies condition $\mathrm{(O)}$. Since $H$ is not normal in $G$, $H/N$ is not normal in $G/N$, and thus by minimality of the counterexample, $H/N$ must be solvable. Now let $x\in G\setminus H$ be an element of odd order. Then $xN\subseteq xH$, and by condition $\mathrm{(O)}$, all elements of $xN$ have odd order. Hence, by [GN, Theorem B], $N$ is solvable. Since $N\trianglelefteq H$ and both $N$ and $H/N$ are solvable, it follows that $H$ is solvable, a contradiction. Therefore, $H_{G}=1$, as wanted.
Claim $6$. There exists a unique maximal subgroup $M$ of $G$ containing $H$. Moreover, $H\trianglelefteq M$ and $M/H$ is a $2$-group. In particular, $H=O^{2}(M)$.
Let $M$ be a maximal subgroup of $G$ containing $H$. If $H=M$, the claim is immediate. Otherwise, suppose $H<M$. Then the pair $(M,H)$ satisfies condition $\mathrm{(O)}$.
By Claim 1, either $H\trianglelefteq M$ and $M/H$ is a $2$-group, or $H$ is solvable. Since $H$ is not solvable, it follows that $H\trianglelefteq M$ and $M/H$ is a $2$-group. Moreover, since $H$ is not normal in $G$, it follows that $M=N_{G}(H),$ which establishes the uniqueness of such a maximal subgroup $M$. Set $K=O^{2}(M)$. Because $H\trianglelefteq M$ and $M/H$ is a $2$-group, we have $K\trianglelefteq H$. But since $H=O^{2}(H)$ and $H/K$ is a $2$-group, this forces $H=K,$ completing the proof.
Claim $7$. If $N$ is a minimal normal subgroup of $G$, then $G=HN.$
Let $N$ be a minimal normal subgroup of $G$. Suppose, for contradiction, that $HN<G$. By Claim 6, $H\leqslant HN\leqslant M$, where $M$ is the unique maximal subgroup containing $H$. It follows that $HN/N$ is a proper subgroup of $G/N$. Thus, either $HN/N\trianglelefteq G/N$ or the quotient $HN/N\cong H/(H\cap N)$ is solvable. (Note that $O^{2}(HN/N)=HN/N.$)
If the former holds, then $H\leqslant HN\trianglelefteq G$, contradicting Claim 4 that the normal closure of $H$ is all of $G$. Hence, the latter must hold, i.e., $H/(H\cap N)$ is solvable. If $N$ is solvable, then $H\cap N$ is also solvable, which would imply $H$ is solvable, a contradiction. Therefore, $N$ is nonsolvable. Since $H_{G}=1$ by Claim 5, we have $N\not\leqslant H$ and consequently $H\cap N<N.$ By Claim 1, either $O^{2}(H\cap N)\trianglelefteq N$ and $N/O^{2}(H\cap N)\text{ is a }2\text{-group},$ or $H\cap N$ is solvable. The latter is impossible, so the former must hold. However, this contradicts Lemma 5.3. Thus, our assumption that $HN<G$ is false, and we conclude that $G=HN$.
Claim $8$. $G$ is a finite nonabelian simple group.
@@
Therefore, $N\cong S^{k}$, where $S$ is a nonabelian simple group and $k\geqslant 1$. Let $U=H\cap N$. Then $U<N$. If every element in $N\setminus U$ has even order, then by Lemma 5.1, $O^{2}(U)\trianglelefteq N$ and $N/O^{2}(U)$ is a nontrivial $2$-group, which is impossible by Lemma 5.3. Thus we can find $x\in N\setminus U$ of odd order. Note that $U\trianglelefteq H$.

### v2

@@ 1. Introduction
Our first theorem establishes the equivalence of the $\mathrm{(CI)}$ and $\mathrm{(F)}$ conditions for any finite group $G$ and nontrivial proper subgroup $H$.
@@
Let $G$ be a finite group and let $H$ be a nontrivial proper subgroup of $G$. Suppose that $H$ is not normal in $G$. Let $N$ be the normal closure of $H$ in $G$. If the pair $(G,H)$ satisfies condition $\mathrm{(F^{\prime})}$, then $(G,N)$ satisfies $\mathrm{(F^{\prime})}$ and $N$ is nilpotent.
@@
Let $G$ be a finite group and let $H$ be a nontrivial proper subgroup of $G$. Study the structure of $G$ if the pair $(G,H)$ satisfies the condition: for every $x\in G\setminus H$, all elements in the coset $xH$ have the same order as $x$.
The main inspiration for our definition of condition $\mathrm{(F)}$ follows from the work of Navarro and Guralnick in [GN]. In that paper, the authors investigate the structure of finite group $G$ with a conjugacy class $K$ such that $K^{2}$ is a conjugacy class of $G$. This is a special case of a conjecture proposed by Arad and Herzog in [AH] stating that in a finite nonabelian simple group, the product of two nontrivial conjugacy classes cannot be a single class. In [GN], they proved that if $K^{2}$ is a conjugacy class of $G$, then $N:=[x,G]$ is solvable, where $x\in K$. Moreover, all elements of the coset $xN$ are conjugate in $G$. Inspired by this, they also show that if $N$ is a normal subgroup of $G$ and $x\in G$ and that either all elements of $xN$ are $G$-conjugate or all elements of $xN$ have odd order, then $N$ is solvable. As before, dropping the normality assumption, we can prove the following result.
@@ 3. $(c)$
This result may be known to experts. In fact, it has been shown that if $x\in G$ and $xy$ is $p$-regular for every $p$-regular element $y\in G$, then $x\in O_{p^{\prime}}(G)$. This was independently proved by Robinson [R] and Flavell [F]. Robinson’s proof is character-theoretic, while Flavell’s argument is group-theoretic. Both proofs do not use the classification of finite simple groups. Our proof of Theorem 5 is also classification-free. Theorem 5 can be considered a dual of Theorem D in [BLM+], where it is shown that if $x\in G$ is $p$-regular such that the order of $xy$ is divisible by $p$ for all nontrivial $p$-elements $y\in G$, then $x\in O_{p^{\prime}}(G)$. This is proved modulo a conjecture on finite almost simple groups.
@@ 2. Preliminaries
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(F)}$ and $G$ is a nilpotent group. Then $H$ is a normal subgroup of $G$.
@@
Since $G$ is nilpotent, it possesses an upper central series
@@
Because $Z_{1}=\mathbf{Z}(G)$ is a nontrivial normal subgroup of $G$, by Lemma 2.6, $Z_{1}\leqslant H$. If $Z_{1}=H$, then the claim follows. Suppose instead that $Z_{1}<H$. Then, by Lemma 2.5, the pair $\left(G/Z_{1},H/Z_{1}\right)$ satisfies condition $\mathrm{(F)}$. Applying Lemma 2.6 to the pair $(G/Z_{1},H/Z_{1})$ yields $Z_{2}/Z_{1}=\mathbf{Z}(G/Z_{1})\leqslant H/Z_{1}$ and hence $Z_{2}\leqslant H$. If $Z_{2}=H$, then we are done. Otherwise, we continue the process: the pair $(G/Z_{2},H/Z_{2})$ satisfies condition $\mathrm{(F)}$ so $Z_{3}/Z_{2}=\mathbf{Z}(G/Z_{2})\leqslant H/Z_{2}$, and thus $Z_{3}\leqslant H$. By repeating this argument, we either eventually reach $H=Z_{i}$ for some $i$, in which case $H\unlhd G$ or else, we find that $Z_{k-1}\leqslant H$. Since $G^{\prime}\leqslant Z_{k-1}\leqslant H$, $H$ is normal in $G$ as well. ∎
@@
Thus, $Kh^{g}=K$ for all $g\in G,h\in H$. Since every $n\in N$ is of the form $h^{g}$ for some $h\in H$ and $g\in G$ by Lemma 2.8, it follows that $Kn=K$ for all $n\in N$, so $KN=K$. Therefore, $xN\subseteq K=x^{G}$ for every $x\in G\setminus N$, and hence $(G,N)$ is a Camina pair. ∎
@@
Since $N$ is nilpotent by Lemma 2.10, its derived subgroup $N^{\prime}$ is a proper characteristic subgroup of $N$. Since $N$ is normal in $G$, $N^{\prime}\unlhd G$. Moreover, since $N$ is the smallest normal subgroup of $G$ containing $H$, we must have $N^{\prime}<H$. Therefore, $H\unlhd N$ and $N/H$ is abelian. ∎
@@ 3. Proof of Theorem 1
where $\theta_{1},\dots,\theta_{t}$ are distinct irreducible characters of $N$ that are conjugate under the action of $G$. Since $\chi_{H},1_{H}]\neq 0$, we may assume without loss of generality that the trivial character $1_{H}$ lies under $\theta_{1},\ldots,\theta_{s}$ for some integer $s$ with $1\leqslant s\leqslant t$. By Lemma 2.4, we know that $H\unlhd N$. Applying Clifford’s Theorem again to each $\theta_{i}$ with $1\leqslant i\leqslant s$, the restrictions of all these characters $\theta_{1},\ldots,\theta_{s}$ to $H$ are multiple of $1_{H}$. Consequently, for $1\leqslant i\leqslant s$, $\theta_{i}$ is an irreducible character of the quotient group $N/H$, which is abelian by Lemma 2.4. Hence these $\theta_{i}$ are all linear characters of $N$. Finally, since $\theta_{1},\dots,\theta_{t}$ are conjugate in $G$, they all have the same degree. Therefore, all $\theta_{i}$, for $1\leqslant i\leqslant t$, are linear characters, as required.
@@
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(CI)}$ and that $H$ is not normal in $G$. Let $N$ be the normal closure of $H$ in $G$. Assume $\chi\in\mathrm{Irr}(G\mid N)$ such that $\chi_{H},1_{H}]=0$. Then $\chi(xh)=\chi(x)=0$ for all $x\in N\setminus H$ and $h\in H$.
@@
where $\theta_{i}$ for $i=1,2,\ldots,t$ are distinct irreducible characters of $N$, which are conjugate in $G$. Since $\chi_{H},1_{H}]=0$, we can find a nontrivial irreducible constituent $\mu_{1}$ of $\chi_{H}$. We may assume that $\theta_{1},\ldots,\theta_{l_{1}}$, for some $l_{1}\leqslant t$, are precisely the irreducible constituents of $\chi_{N}$ lying over $\mu_{1}$. Analogous to the argument in Proposition 3.1, we claim
@@
Let $N$ denote the normal closure of $H$ in $G$. By Lemma 2.2, $(G,N)$ is a Camina pair. That is, for every $\in G\setminus N, xx$ is conjugate to $xn$ for all $n\in N$. In particular, for all $x\in G\setminus N$ and $h\in H\subseteq N$, $x$ is conjugate to $xh$, as required for condition $\mathrm{(F)}$. It remains to consider the case when $x\in N\setminus H$. We claim that for all $\in N\setminus H, xx$ is conjugate to $xh$ for all $h\in H$. This is equivalent to the fact that, for all $x\in N\setminus H$, $\chi(x)=\chi(xh)$ for all $h\in H$ and $\chi\in\mathrm{Irr}(G)$. Indeed, for any $\chi\in\mathrm{Irr}(G/N)$, we have $\chi(x)=\chi(xh)=\chi(1)$ for all $x\in N,h\in H$ as $N$ is contained in the kernel of $\chi$. For $\chi\in\mathrm{Irr}(G\mid N)$, the equality $\chi(x)=\chi(xh)$ for all $x\in N\setminus H$, $h\in H$ follows from Propositions 3.1 and 3.2. Thus, all irreducible characters of $G$ are constant on the cosets $xH$ for $x\in N\setminus H$, so $xH\subseteq x^{G}$ for all $x\in N\setminus H.$ Hence $(G,H)$ satisfies condition $\mathrm{(F)}$, completing the proof. ∎
@@
If $H$ is normal in $G$, then this is Lemma 2.1 in [KV]. Assume that $H$ is not normal in $G$. Then the result follows from Propositions 3.3 and 3.4. ∎
@@ 4. Proof of Theorem 2
Let $G$ be a finite group and let $H$ be a subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(F^{\prime})}$ and that $H$ is not normal in $G$. Let $N$ be the normal closure of $H$ in $G$. Then $N$ is solvable and has a normal $p$-complement for some prime $p$ dividing $|G|$.
@@
Since $H$ is a proper subgroup of $G$, $\Delta_{H}(G)=G\setminus\cup_{g\in G}H^{g}$ is non-empty. By Corollary 4.4, we have $\Delta_{H}(G)=G\setminus N$ and so $N<G$. Let $x\in\Delta_{H}(G)$ and let $K=x^{G}$. By Lemma 4.3, $KN\subseteq K\cup K^{-1}$, and hence $xN\subseteq K\cup K^{-1}$. By [B, Theorem 3.3(a)], NN is solvable. Since $G/N$ is nontrivial, we can find $x\in G\setminus N$ such that $x$ is a $p$-element for some prime $p$. By [B, Theorem 3.3(c)], NN has a normal $p$-complement. ∎
@@
We now prove that $G$ is a Frobenius group with Frobenius kernel $N$, so that $N$ is nilpotent by Thompson’s Theorem, yielding a contradiction. Recall that $N$ is a non-nilpotent normal $p$-complement of $G$. Since $G=PN$, with $N$ a normal $p$-complement and $P\cap N=1$, it suffices (see [I, Problem 7.1]) to show that $\mathbf{C}_{G}(n)\subseteq N$ for every $1\neq n\in N.$ Suppose, for contradiction, that there exists $x\in\mathbf{C}_{G}(n)\setminus N$ for some $1\neq n\in N$. Then $x$ is a $p$-element and $xn=nx$. Since $n$ is a $p^{\prime}$-element, the order of $xn$ is $o(x)o(n)$, so $xn$ is not a $p$-element. But since $x\in G\setminus N$, condition $\mathrm{(F^{\prime})}$ implies that $xn\in xN\subseteq x^{G}\cup(x^{-1})^{G}$, a set consisting of only $p$-elements. This contradiction shows that $\mathbf{C}_{G}(n)\subseteq N$, and thus $G$ is a Frobenius group with Frobenius kernel $N$. By Thompson’s theorem [Th, Theorem 1], $N$ is nilpotent, contradicting our assumption. Therefore, $N$ must be nilpotent, completing the proof. ∎
@@ 2. $(2)$
Let $K_{0}$ be the subgroup of $G$ generated by all $p$-regular elements of $G$. Then $K_{0}\unlhd G$ because the set of $p$-regular elements is closed under conjugation. We will show that $K_{0}=K$, and so $K\unlhd G$. From the previous paragraph, $K_{0}\subseteq K$. Now consider an arbitrary element $hK_{0}\in H/K_{0}$, and write $h=h_{p}h_{p^{\prime}}$ where $h_{p},h_{p^{\prime}}$ are the $p$-part and $p^{\prime}$-part of $h$, respectively. Then $h_{p^{\prime}}\in K_{0}$, so $hK_{0}=h_{p}K_{0}$ has $p$-power order in $H/K_{0}$. Thus all elements of $H/K_{0}$ are $p$-elements, and so $H/K_{0}$ is a $p$-group. Hence $K\subseteq K_{0}$, and we conclude that $K=K_{0}\unlhd G$. Moreover, $|G:H|$ is a $p$-power and $H/K$ is a $p$-group, which implies that $G/K$ is a $p$-group.
@@
We make the following remark.
Remark 5.3.
@@
Let $G$ be a finite group and let $H$ be a proper subgroup of $G$. Suppose that the pair $(G,H)$ satisfies condition $\mathrm{(O)}$. Then either $O^{2}(H)\unlhd G$ or $H$ is solvable.
@@
This claim follows immediately from condition $\mathrm{(O)}$. If $gH$ has an element of odd order, then all elements in $gH$ also have odd order, and so is $g$. However, $g\in G\setminus H$ is assumed to have even order, a contradiction.
@@
Suppose, for contradiction, that $N=H_{G}>1$. Then $N\trianglelefteq G$, $N\leqslant H$, and $1<N<H$. By Claim 3, the pair $(G/N,H/N)$ satisfies condition $\mathrm{(O)}$. Since $H$ is not normal in $G$, $H/N$ is not normal in $G/N$, and thus by minimality of the counterexample, $H/N$ must be solvable. Now, as $H=O^{2}(H)$ is not normal in $G$, by Lemma 5.1, there exists $x\in G\setminus H$ of odd order. Then $xN\subseteq xH$, and by condition $\mathrm{(O)}$, all elements of $xN$ have odd order. Hence, by [GN, Theorem B], $N$ is solvable. Since $N\trianglelefteq H$ and both $N$ and $H/N$ are solvable, it follows that $H$ is solvable, a contradiction. Therefore, $H_{G}=1$, as wanted.
Claim $6$. If $N$ is a minimal normal subgroup of $G$, then $G=HN.$
Let $N$ be a minimal normal subgroup of $G$. Suppose, for contradiction, that $HN<G$. It follows that $HN/N$ is a proper subgroup of $G/N$. Thus, either $HN/N\trianglelefteq G/N$ or the quotient $HN/N\cong H/(H\cap N)$ is solvable. (Note that $O^{2}(HN/N)=HN/N.$)
If the former holds, then $H\leqslant HN\trianglelefteq G$, contradicting Claim 4 that the normal closure of $H$ is all of $G$. Hence, the latter must hold, i.e., $H/(H\cap N)$ is solvable. If $N$ is solvable, then $H\cap N$ is also solvable, which would imply $H$ is solvable, a contradiction. Therefore, $N$ is nonsolvable. Since $H_{G}=1$ by Claim 5, we have $N\not\leqslant H$ and consequently $H\cap N<N.$ By Claim 1, either $O^{2}(H\cap N)\trianglelefteq N$ and $N/O^{2}(H\cap N)\text{ is a }2\text{-group},$ or $H\cap N$ is solvable. The latter is impossible, so the former must hold. However, this contradicts Remark 5.3. Thus, our assumption that $HN<G$ is false, and we conclude that $G=HN$.
Claim $7$. $G$ is a finite nonabelian simple group.
@@
Therefore, $N\cong S^{k}$, where $S$ is a nonabelian simple group and $k\geqslant 1$. Let $U=H\cap N$. Then $U<N$. If every element in $N\setminus U$ has even order, then by Lemma 5.1, $O^{2}(U)\trianglelefteq N$ and $N/O^{2}(U)$ is a nontrivial $2$-group, which is impossible by Remark 5.3. Thus we can find $x\in N\setminus U$ of odd order. Note that $U\trianglelefteq H$.